# Context Parallelism for Scalable Million-Token Inference

### v2

@@ Abstract
[…]
Jianyu Huang at
@@ 1 Introduction
Contemporary large language models (LLMs), such as Llama Touvron et al. (2023a, b); Llama Team (2024), Gemini Gemini Team (2023, 2024), GPT-4 Achiam et al. (2023), require significant computational resources for inference, especially with long context lengths: OpenAI GPT-4o 128K context length ope , Anthropic’s Claude with 200K context length ant , Google’s Gemini 1.5 Pro with 1M context length goo . With a single H100 GPU host (8 GPUs), it can take 60 seconds to serve 128K context length [^1] Google’s Gemini 1.5 Pro (Sep 2024) has a latency of 20.43 seconds for time to first token on 100K context length, from https://artificialanalysis.ai/models/gemini-1-5-pro/prompt-options/single/100k. or 1200 seconds to serve 1M context length for Llama3 405B model. Context parallelism (CP) is a system optimization technique that improves the latency and scalability of LLM inference, particularly for long contexts. Without modifying the underlying dense attention algorithms, CP offers several advantages for long-context LLM inference:
@@ 2.3 Prior works on Long Context LLM
  • System-level optimizations: preserve the model architecture, instead improve the scalability of existing dense attention algorithms to leverage more compute resources Li et al. (2021); Brandon et al. (2023); Liu et al. (2023).
@@ 3.3 Computation and Communication Modeling
To understand whether communication can be reliably overlapped with attention computation with varying persistent KV hit rates, we approximate the attention computation and QKV communication latency using a simple roof-line model.
@@
Note that the threshold for $T$, the length of new tokens is a static threshold with respect to a given model and hardware, which is independent of KV cache hit $P$.
@@
To summarize, we adaptively switch between pass-KV and pass-Q for inference partial prefill following the heuristics in Algorithm 1222In practice, the achieved BW and C are lower than the theoretical hardware peaks. We start with these peak values and then fine-tune the thresholds based on empirical data.. We can calculate the static thresholds for this heuristics once based on the system and model spec, and use the heuristics to choose which options to use dynamically for the optimal performance in a wide combination of total context length and KV cache hit thresholds.
@@ 3.4.2 Ring Pass-KV Algorithm
At the end of the ring algorithm loop, each CP rank $k$ will have the attention output of $O_{k}^{s}$ with $s=0,1,...,N-1$, where $O_{k}^{s}$ denotes the attention output from $Q_{k}$ and $KV^{s}$ (key and value embeddings originally sharded to rank $s$, see bottom of Figure 3). We then apply a merge attention operator Juravsky et al. (2024) to get the result of $Q_{k}$ interacted with all $KV$ embeddings across CP ranks (See Appendix C, Equation (4)).
@@ 3.4.3 Ring Pass-Q Algorithm
All2All for partial attention outputs is on the critical path and therefore introduces an additional communication overhead apart from the communication for passing query embedding. The analysis for overlapping query embedding and attention in Equation (2) and (3) only applies to the ring communication. The heuristics in Algorithm 1 for switching between pass-KV and pass-Q doesn’t take All2All latency into account [^3] We present a refined algorithm in Appendix D and provide a detailed time breakdown for validations in Table 4..
@@ 4.2.1 Latency Reduction with Fixed Context Length
(b) GTI Latency for CP with 1, 2, 4 nodes.
@@ 4.2.3 Scaling Context Length with Fixed Capacity
We calculate the FLOPS utilization of a 1M context length on 16 nodes in Appendix B. The achieved FLOPS is 502 TF/sec per H100, compared to a standalone Flash Attention v3 benchmark performance of 540 TF/sec for 8K context length (1M over 128 H100 GPUs) on a single GPU, resulting in a 93% parallelization efficiency. Considering the peak FLOPS on the specialized H100 configurations, we achieve approximately 63% FLOPS utilization.
@@
Table 3: TTFT (in $ms$) for pass-KV vs. pass-Q varying $P$ and $T$ with $P+T=128000$, on 4 CP ranks (CP4).
@@
We further validate the analytical model in Algorithm 1 for predicting the selection of pass-KV vs. pass-Q from Table 3. When the KV cache miss rate exceeds 12.5% ($2\cdot\frac{N_{KV}}{N_{H}}$ in Equation 1), pass-KV is always selected, meeting the 2nd condition in Algorithm 1. At 10% KV cache miss rate, pass-KV remains the choice since the number of new tokens $T$ is sufficiently large, satisfying Equation 2 (with SendRecv hidden under Attn in Table 4); Around 5% cache miss rate (e.g., $T=6400$), the differences between pass-KV and pass-Q is less than 1%, allowing for either option to be selected. When cache miss rate falls below 3.25%, pass-KV communication becomes exposed, leading to the selection of pass-Q. Specifically, at a 2.5% cache miss rate, the sum of the exposed communication in pass-KV ring loop is less than All2All exposed in pass-Q, resulting in the selection of pass-Q .
@@ 4.3 Decode Performance
Parallelism Scalability: We benchmarked different parallelization configurations up to four CP nodes to observe the scalability of both TP and CP. Table 6 shows that TTIT tends to be longer for both scaling TP and scaling CP. TTIT for scaling TP increases to 47 $ms$ while TTIT for scaling CP increases to 71 $ms$. Both TP and CP have poor scalability for decoding because when adding more hosts. For TP, lower computation latency on linear layers is offset by increased communication latency increased.
For CP, as we increase the number of hosts, the effective length seen by each attention kernel decreases, so each individual attention op becomes faster (Table  7). However TTIT still degrates compared to CP=1, and the reason for that is two-fold: 1) Current implementation pads the number of queries to make it divisible by the number of ranks, which for B=1 means the total number of processed queries increases with CP. 2) The communication latency - sending Q chunks to the next rank at each iteration of the loop and all2all-exchanging partial attention outputs after the loop - also grows with the number of hosts. As a result, the total pass-Q attention latency and TTIT increase with CP.
@@ 6 Acknowledgments
Appendix A Notations
We attach the notations used in this paper for reference in Table 8.
Table 8: Notation table.
         Notation          Description
         $N_{H}$          Number of query heads (or attention heads)
         $N_{KV}$          Number of key/value heads
         $D_{H}$          Head dimension in Transformer
         $D$          Model dimension in Transformer: $D_{H}\cdot N_{H}$
         $Q,K,V$          Query, Key, Value tensors
         $B$          Batch size (# of input sequences)
         $W$          Model parameter size
         $T$          Input sequence length
         $P$          Previously cached KV length
         $e$          Element data type size
         $C$          Peak compute
         $BW$          Peak comm bandwidth
         $bid$          Decode batch id
          $T^{i}_{j}|P^{i}_{j}$          # of {new tokens |||| cached KV tokens} in $i$-th
         sequence sharded to CP rank $j$
         $L^{i}$          $KV$ length in $i$-th sequence: $max_{j=0}^{N-1}(P^{i}_{j}+T^{i}_{j})$
         $N$          Number of hosts/nodes/CP ranks
         $N_{TP}$          TP group size (# of GPUs in a TP group)
          $Q^{s}_{k}|KV^{s}_{k}|bid^{s}_{k}$          {Query |||| key and value tensors |||| batch id }
         on rank $k$, originally allocated to rank $s$
         $O^{s}_{k}$          Attention output from $Q_{k}$ and $KV^{s}$
         $TP8,TP16$          Tensor parallel sharding over 8 GPUs
         on one node or 16 GPUs on two nodes
          $CP_{N}$          Context parallel sharding on $N$ nodes
         with TP8 for each node (same as $CP_{N}$+$TP8$)
          $TTFT$          Time-to-first-token: latency for prefilling
         the whole input tokens
          $TTIT$          Time-to-incremental-token: latency for
         decoding each output token
Appendix B MFU Calculation for 1M context length
@@
Appendix C Merge Attention
The idea of merging attention outputs from different keys/values originates from Online Softmax Milakov and Gimelshein (2018). Later this idea was reused in Flash Attention Dao et al. (2022); Dao (2023). Here we derive the equation to merge the partial attention outputs from different CP ranks.
@@ Appendix C Merge Attention
$O=\sum_{i=0}^{R-1}V_{i}\cdot e^{Q\cdot K_{i}^{T}/\sqrt{d}}e^{-LSE},$
@@
$\large O_{k}=\frac{\sum_{s=0}^{N-1}O_{k}^{s}\times e^{LSE_{k}^{s}-LSE_{k}^{max}}}{\sum_{i}^{N-1}e^{LSE_{k}^{s}-LSE_{k}^{max}}},$ (4)
@@
Appendix D Analytical Model Selection Considering All2All
pass-Q merge attention requires an All2All (Section 3.4.3), whereas in pass-KV merge attention only needs to merge the partial attn results on local node (Section 3.4.2). When pass-KV communication is exposed, we want to compare the total of exposed pass-KV’s communication time to the pass-Q’s all2all, which is the time to send partial attention output and partial attention softmax log-sum-exp (LSE) (Appendix C):
@@ Appendix D Analytical Model Selection Considering All2All
$\geq(N-1)\cdot\frac{(D+1)\cdot T\cdot e}{BW}$
@@
Appendix E Heuristic based on empirical data

### v3

@@ Abstract
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Contemporary large language models (LLMs), such as Llama Touvron et al. (2023a; b); Llama Team (2024), Gemini Gemini Team (2023; 2024), GPT-4 Achiam et al. (2023), require significant computational resources for inference, especially with long context lengths: OpenAI GPT-4o 128K context length ope , Anthropic’s Claude with 200K context length ant , Google’s Gemini 1.5 Pro with 1M context length goo . With a single H100 GPU host (8 GPUs), it can take 60 seconds to serve 128K context length [^1] Google’s Gemini 1.5 Pro (Sep 2024) has a latency of 20.43 seconds for time to first token on 100K context length, from https://artificialanalysis.ai/models/gemini-1-5-pro/prompt-options/single/100k. or 1200 seconds to serve 1M context length for Llama3 405B model. Context parallelism (CP) is a system optimization technique that improves the latency and scalability of LLM inference, particularly for long contexts. Without modifying the underlying dense attention algorithms, CP offers several advantages for long-context LLM inference:
@@ 2.3 Prior works on Long Context LLM
  • System-level optimizations: preserve the model architecture, instead improve the scalability of existing dense attention algorithms to leverage more compute resources Li et al. (2021); Brandon et al. (2023); Liu et al. (2023); Wu et al. (2024); Li et al. (2023); Jacobs et al. (2023); Fang & Zhao (2024).
@@ 3 Context Parallel Inference
3.1 Notations
The notations used in this paper are summarized in Table 1.
Table 1: Notation table.
Notation Description
$N_{H}$ Number of query heads (or attention heads)
$N_{KV}$ Number of key/value heads
$D_{H}$ Head dimension in Transformer
$D$ Model dimension in Transformer: $D_{H}\cdot N_{H}$
$Q,K,V$ Query, Key, Value tensors
$B$ Batch size (# of input sequences)
$W$ Model parameter size
$T$ Input sequence length
$P$ Previously cached KV length
$e$ Element data type size
$C$ Peak compute
$BW$ Peak comm bandwidth
$bid$ Decode batch id
$T^{i}_{j}|P^{i}_{j}$ of {new tokens |||| cached KV tokens} in $i$-th
sequence sharded to CP rank $j$
$L^{i}$ $KV$ length in $i$-th sequence: $max_{j=0}^{N-1}(P^{i}_{j}+T^{i}_{j})$
$N$ Number of hosts/nodes/CP ranks
$N_{TP}$ TP group size (# of GPUs in a TP group)
$Q^{s}_{k}|KV^{s}_{k}|bid^{s}_{k}$Query |||| key and value tensors |||| batch id }
on rank $k$, originally allocated to rank $s$
$O^{s}_{k}$ Attention output from $Q_{k}$ and $KV^{s}$
$TP8,TP16$ Tensor parallel sharding over 8 GPUs
on one node or 16 GPUs on two nodes
$CP_{N}$ Context parallel sharding on $N$ nodes
with TP8 for each node (same as $CP_{N}$+$TP8$)
$TTFT$ Time-to-first-token: latency for prefilling
the whole input tokens
$TTIT$ Time-to-incremental-token: latency for
decoding each output token
@@ 3.4 Computation and Communication Modeling
To understand whether communication can be reliably overlapped with attention computation with varying persistent KV hit rates, we approximate the attention computation and QKV communication latency using a simple roof-line model (Table 3).
@@
Note that the threshold for $T$, the length of new tokens is a static threshold with respect to a given model and hardware, which is independent of KV cache hit on the previously cached KV length $P$.
@@
To summarize, we adaptively switch between pass-KV and pass-Q for inference partial prefill following the heuristics in Algorithm 1222In practice, the achieved BW and C are lower than the theoretical hardware peaks. We start with these peak values and then fine-tune the thresholds based on empirical data.. It’s worth noting that the full prefill can be considered as a special case where $P=0$, while decoding can be viewed as a special case where $T=1$. We can calculate the static thresholds for this heuristics once based on the system and model spec, and use the heuristics to choose which options to use dynamically for the optimal performance in a wide combination of total context length and KV cache hit thresholds.
@@ 3.5.2 Ring Pass-KV Algorithm
At the end of the ring algorithm loop, each CP rank $k$ will have the attention output of $O_{k}^{s}$ with $s=0,1,...,N-1$, where $O_{k}^{s}$ denotes the attention output from $Q_{k}$ and $KV^{s}$ (key and value embeddings originally sharded to rank $s$, see bottom of Figure 3). We then apply a merge attention operator Juravsky et al. (2024) to get the result of $Q_{k}$ interacted with all $KV$ embeddings across CP ranks (See Appendix B, Equation (4)).
@@ 3.5.3 Ring Pass-Q Algorithm
All2All for partial attention outputs is on the critical path and therefore introduces an additional communication overhead apart from the communication for passing query embedding. The analysis for overlapping query embedding and attention in Equation (2) and (3) only applies to the ring communication. The heuristics in Algorithm 1 for switching between pass-KV and pass-Q doesn’t take All2All latency into account [^3] We present a refined algorithm in Appendix C and provide a detailed time breakdown for validations in Table 5..
@@ 4.2.1 Latency Reduction with Fixed Context Length
(b) GTI Latency for CP with 1, 2, 4 nodes
@@ 4.2.3 Scaling Context Length with Fixed Capacity
We calculate the FLOPS utilization of a 1M context length on 16 nodes in Appendix A. The achieved FLOPS is 502 TF/sec per H100, compared to a standalone Flash Attention v3 benchmark performance of 540 TF/sec for 8K context length (1M over 128 H100 GPUs) on a single GPU, resulting in a 93% parallelization efficiency. Considering the peak FLOPS on the specialized H100 configurations, we achieve approximately 63% FLOPS utilization.
@@
Table 4: TTFT (in $ms$) for pass-KV vs. pass-Q varying $P$ and $T$ with $P+T=128000$, on 4 CP ranks (CP4). $P$: length of existing tokens in the KV cache, $T$: length of new tokens.
@@
We further validate the analytical model in Algorithm 1 for predicting the selection of pass-KV vs. pass-Q from Table 4.
  • When the KV cache miss rate exceeds 12.5% ($=2\cdot\frac{N_{KV}}{N_{H}}$ in Equation 1), pass-KV is always selected, meeting the 2nd condition in Algorithm 1.
  • At 10% KV cache miss rate, pass-KV remains the choice since the number of new tokens $T$ is sufficiently large, satisfying Equation 2 (with SendRecv hidden under Attn in Table 5).
  • Around 5% cache miss rate (e.g., $T=6400$), the differences between pass-KV and pass-Q is less than 1%, allowing for either option to be selected.
  • When cache miss rate falls below 3.25%, pass-KV communication becomes exposed, leading to the selection of pass-Q. Specifically, at a 2.5% cache miss rate, the sum of the exposed communication in pass-KV ring loop is larger than All2All exposed in pass-Q (Equation 5, Appendix C), resulting in the selection of pass-Q .
@@ 4.3 Decode Performance
Parallelism Scalability: We benchmarked different parallelization configurations up to four CP nodes to observe the scalability of both TP and CP. Table 7 shows that TTIT tends to be longer for both scaling TP and scaling CP. TTIT for scaling TP increases to 47 $ms$ while TTIT for scaling CP increases to 71 $ms$. Both TP and CP have poor scalability for decoding when adding more hosts (e.g., using 4 nodes can result in worse TTIT than using a single node). For TP, lower computation latency on linear layers is offset by increased communication latency increased.
For CP, as we increase the number of hosts, the effective length seen by each attention kernel decreases, so each individual attention op becomes faster (Table  8). However TTIT still degrates compared to CP=1, and the reason for that is two-fold: (1) Current implementation pads the number of queries to make it divisible by the number of ranks, which for B=1 means the total number of processed queries increases with CP. (2) The communication latency - sending Q chunks to the next rank at each iteration of the loop and All2All-exchanging partial attention outputs after the loop - also grows with the number of hosts. As a result, the total pass-Q attention latency and TTIT increase with CP.
@@ 6 Acknowledgments
Appendix A MFU Calculation for 1M context length
@@
Appendix B Merge Attention
The idea of merging attention outputs from different keys/values originates from Online Softmax Milakov & Gimelshein (2018). Later this idea was reused in Flash Attention Dao et al. (2022); Dao (2023). Here we derive the equation to merge the partial attention outputs from different CP ranks.
@@ Appendix B Merge Attention
$O=\frac{\sum_{i=0}^{R-1}\text{\tt exp}^{Q\cdot K_{i}^{T}/\sqrt{d}}\cdot V_{i}}{\text{\tt exp}^{LSE}},$
@@
$O_{k}=\frac{\sum_{s=0}^{N-1}(O_{k}^{s}\times\text{\tt exp}^{LSE_{k}^{s}-LSE_{k}^{max}})}{\sum_{s=0}^{N-1}\text{\tt exp}^{LSE_{k}^{s}-LSE_{k}^{max}}},$ (4)
@@
Appendix C Analytical Model Selection Considering All2All
pass-Q merge attention requires an All2All (Section 3.5.3), whereas in pass-KV merge attention only needs to merge the partial attn results on local node (Section 3.5.2). When pass-KV communication is exposed, we want to compare the total of exposed pass-KV’s communication time to the pass-Q’s all2all, which is the time to send partial attention output and partial attention softmax log-sum-exp (LSE) (Appendix B):
@@ Appendix C Analytical Model Selection Considering All2All
$\geq(N-1)\cdot\frac{(D+1)\cdot T\cdot e}{BW}.$
@@
Appendix D Heuristic based on empirical data